# Variation of the centered maximal operator

### v2

@@ 1. Introduction
It was an important observation of Aldaz and Pérez Lázaro [1] that maximal operators can actually improve the regularity of involved functions rather than simply to preserve it. They also pointed out the role to be played by variation when measuring the regularity. In particular, it was proved in [1] that if $f\in{\rm BV}(\mathbb{R})$, then for the uncentered Hardy–Littlewood maximal operator $\widetilde{M}$ the function $\widetilde{M}f$ is absolutely continuous on $\mathbb{R}$ and ${\rm Var}(\widetilde{M}f)\leq{\rm Var}(f)$. For the centered operator $M$, on the other hand, Kurka [12] proved
@@
with some, quite large, $C>0$. It was conjectured that also in this case $C=1$ is enough, see e.g. [12, Section 1]. In the discrete case, i.e., for $f\in{\rm BV}(\mathbb{Z})$ and the discrete counterpart of $M$, the same was conjectured, see [3, Question B]. It is worth recalling the common truth that the uncentered maximal operator has better regularity behavior than its centered counterpart. Furthermore, the proofs of analogous results are usually much subtle in the centered setting.
It seems that both conjectures, although reasonable and expected to be true, do not take the full advantage of how the limits $f(\pm\infty)$ or $F(\pm\infty)$ for $f\in{\rm BV}(\mathbb{R})$ or $F\in{\rm BV}(\mathbb{Z})$, respectively, determine the limits $Mf(\pm\infty)$ or $MF(\pm\infty)$; see Section 3. We thus propose the following strengthening.
@@
Our first main result, Theorem 1.1, supports this conjecture. We mention that there is no need to formulate separately the discrete counterpart of Conjecture $\Delta$, since by our second main result, Theorem 2.2, inequality (1.1) and its discrete analogue are equivalent. Clearly, verification of either version of the conjecture, continuous or discrete, reduces to consideration of nonnegative functions. We shall prove that a certain further reduction to nonnegative simple functions is available; see Proposition 2.3. Finally, we show that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal; see Section 4.
Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{Z}$. For $g\colon\mathbb{K}\to\mathbb{C}$ and any nonempty $E\subseteq\mathbb{K}$ define
@@
for any locally integrable $g\colon\mathbb{R}\to\mathbb{C}$ and $x\in\mathbb{R}$, and for any $G\colon\mathbb{Z}\to\mathbb{C}$ and $n\in\mathbb{Z}$. Of course, if $g\in{\rm BV}(\mathbb{R})$, then $g$ is bounded, hence so is $Mg$. The same is true for $MG$ when $G\in{\rm BV}(\mathbb{Z})$.
@@
Let $-\infty<A<B<+\infty$, $a,b\in\mathbb{C}$, $K\in\mathbb{N}\coloneqq\{1,2,\dots\}$, and $\{\alpha_{k}\}_{1}^{K}\subset\mathbb{C}$ be given. For a system $-\infty<A=t_{0}<t_{1}<\dots<t_{K}=B<+\infty$ define
$f(x)\coloneqq a{\mathds{1}_{{(-\infty,A)}}}(x)+\sum_{k=1}^{K}\alpha_{k}{\mathds{1}_{{(t_{k-1},t_{k})}}}(x)+b{\mathds{1}_{{(B,\infty)}}}(x),$
for $x\in\mathbb{R}\setminus\{t_{0},\ldots,t_{K}\}$. Arbitrary values can be set for $f(t_{0})$ and $f(t_{K})$, and also for $f(t_{k})$, $11\leq k\leq K-$ 1, unless both $\alpha_{k}$ and $\alpha_{k+1}$ are different from 00 when we set $f(t_{k})=0$. Then
@@
Theorem 1.1 implies (1.2) or (1.3) for all indicator functions of bounded variation, $f\colon\mathbb{R}\to\{0,1\}$ or $F\colon\mathbb{Z}\to\{0,1\}$, respectively, thus strengthening [4, Theorems 1.1 and 1.3]. Notice a separation condition encoded in the considered systems $\{t_{k}\}_{0}^{K}$ and $\{n_{k}\}_{0}^{K}$, when we require either $f(t_{k})=0$, in some circumstances depending on the behavior of $f$ around $t_{k}$, or $\alpha_{k}\alpha_{k+1}=0$, respectively.
Throughout the paper, $I$ with possible subscripts affixed will denote an open bounded interval on $\mathbb{R}$. Then $c(I),l(I),r(I)$ will be the center and the left/right endpoint of $I$, respectively. As already mentioned, ${\rm Var}$ with no subscripts, $\mathbb{R}$ or $\mathbb{Z}$, will denote the total variation; it will be clear from the context to which situation this symbol refers to. Appearance of a subscript will indicate that the variation is related to a subset represented by this subscript.
@@ 2. Transference
$M_{<\delta}\tilde{g}(y)=\max\Big{\{}g(x^{+}),\frac{(\delta-y+x)g(x^{-})+(\delta+y-x)g(x^{+})}{2\delta}\Big{\}},$
@@
$|M_{\geq\delta}\tilde{g}(z)-M_{\geq\delta}\tilde{g}(y)|\leq\frac{(z-y){\rm Var}(g)}{2\delta},$
@@
Thus, $\limsup_{t\to x^{+}}Mg(t)-\liminf_{t\to x^{+}}Mg(t)\leq 2\varepsilon$. Analogously we show that $Mg(x^{-})$ exists.
Regarding (2.1), we have $Mg(x)\leq\min\{Mg(x^{-}),Mg(x^{+})\}$ by lower semicontinuity of $Mg$. By using $Mg(x)\geq\frac{g(x^{-})+g(x^{+})}{2}$ and $Mg(x^{-})\geq\max\{g(x^{-}),\frac{g(x^{-})+g(x^{+})}{2}\},$ and standard symmetry arguments, we conclude that to show the opposite inequality it suffices to prove the implication
$Mg(x^{-})>\max\Big{\{}g(x^{-}),\frac{g(x^{-})+g(x^{+})}{2}\Big{\}}\implies Mg(x)\geq Mg(x^{-}).$
But if the left part above holds, then there is $\delta>0$ such that $Mg(y)=M_{\geq\delta}g(y)$ for $y\in(x-\delta,x)$. Since like before $|M_{\geq\delta}g(x)-M_{\geq\delta}g(y)|\leq\frac{(x-y){\rm Var}(g)}{2\delta}$ for all such $y$, the proof is concluded. ∎
@@
Then ${\rm Var}(g)={\rm Var}(G)$ so that $g\in{\rm BV}_{a}^{b}(\mathbb{R})$. Moreover, $Mg(n)=MG(n)$ for $n\in\mathbb{Z}$, since in this case the averages of $|g|$ over the centered intervals $I$ of length $|I|\in 2\mathbb{N}-1$ dominate the remaining ones. Thus, ${\rm Var}(MG)={\rm Var}_{\mathbb{Z}}(Mg)\leq{\rm Var}(Mg)$ and the claimed inequalities follow.
In the opposite direction, we claim that $c(\mathbb{Z})\geq c(\mathbb{R})$ and $C_{a}^{b}(\mathbb{Z})\leq C_{a}^{b}(\mathbb{R})$, and use a convenient variant of a sampling procedure. Obviously, we can assume that $c(\mathbb{Z})<\infty$. Given $g\in{\rm BV}_{a}^{b}(\mathbb{R})$ and $\varepsilon>0$, in view of (2.1), there is $N_{*}=N_{*}(\varepsilon)\in\mathbb{N}$ such that for $E_{N_{*}}\coloneqq 2^{-N_{*}}\mathbb{Z}\cap[-2^{N_{*}},2^{N_{*}}]$ we have
@@
so that $\tilde{G}_{N}\in{\rm BV}_{|a|}^{|b|}(\mathbb{Z})$ and $G_{N}\in{\rm BV}_{a}^{b}(\mathbb{Z})$. Notice that $MG_{N}(n)\leq M\tilde{G}_{N}(n)=M_{N}g(n/2^{N})$ for all $n\in\mathbb{N}$. Now, we show that taking $N\geq\max\{N_{*},N_{0}\}$ large enough we obtain $MG_{N}(n)\geq M\tilde{G}_{N}(n)-\varepsilon/(10N_{*}^{2})$ for all $n\in E_{N_{*}}^{N}\coloneqq\{2^{N}Q:Q\in E_{N_{*}}\}\subset\mathbb{N}$. Given $Q\in E_{N_{*}}$, we have
$|I_{Q}|^{-1}\int_{I_{Q}}|g|\geq M\tilde{G}_{N}(2^{N}Q)-\varepsilon/(20N_{*}^{2})$
for some $I_{Q}$ centered at $Q$. For each $N$ large enough, let $I_{Q}^{N}$ be the longest interval centered at $Q$, contained in $I_{Q}$, with $|I_{Q}^{N}|\in(2\mathbb{N}-1)/2^{N}$. Denoting $g_{N}\coloneqq\sum_{n\in\mathbb{Z}}G_{N}(n){\mathds{1}_{{[(n-1/2)/2^{N},(n+1/2)/2^{N})}}}$, we see that $\lim_{N\to\infty}|I_{Q}^{N}|^{-1}\int_{I_{Q}^{N}}|g_{N}|=|I_{Q}|^{-1}\int_{I_{Q}}|g|$ by the dominated convergence theorem on $I_{Q}$, because $g$ is continuous almost everywhere and bounded. Since $MG_{N}(2^{N}Q)\geq|I_{Q}^{N}|^{-1}\int_{I_{Q}^{N}}|g_{N}|$ and $E_{N_{*}}$ is finite, the desired bounds follow for some large $N\geq\max\{N_{*},N_{0}\}$. Consequently,
@@
which contradicts our assumption $c(\mathbb{Z})<\infty$. Hence, we have ${\rm Var}(Mg)<\infty$ (this includes the case $c(\mathbb{R})<\infty$), and applying successively the first part of (2.2), (2.4), (2.3) for $G_{N}$, and the definitions of $c(\mathbb{Z})$ and $C_{a}^{b}(\mathbb{Z})$, and then letting $\varepsilon\to 0$, we verify the claim. ∎
@@
$\max\{Mg(y^{\prime}),Mg(y^{\prime\prime})\}<\sup\{Mg(y):y\in\mathbb{K}\cap[y^{\prime},y^{\prime\prime}]\}=Mg(x);$
@@
Until the end of this section we fix $f$ as in Theorem 1.1 and denote $I_{k}=(t_{k-1},t_{k})$, $k=1,\ldots,K$, the corresponding intervals. Without any loss of generality we can consider nonnegative $f$ only, hence from now on we assume $a,b\geq 0$ and $\alpha_{k}\geq 0$ for $k=1,\ldots,K$. Clearly, $Mf$ may not have local maxima at all. However, if $x$ is a local maximum of $Mf$, then two cases may occur:
@@
    $Mf$ is constant on some interval $J\ni x$, i.e., each point of $J$ is also a local maximum of $Mf$;
@@
In the first case, considering $J$ as the longest interval with the declared property, we shall choose exactly one element $x_{*}\in J$ as a representative and this local maximum will be called substantial. In addition every isolated maximum from the second case, if any, will also be called substantial. It follows from Lemma 3.1 that the number of substantial maxima is finite and does not exceed $K^{2}$.
@@
Let $x$ be a local maximum of $Mf$. Then there exists $x_{*}$ (either equal to $x$ or such that $Mf$ is constant on the closed interval connecting $x$ and $x_{*}$) with the property that $Mf(x_{*})$ is attained for some $I_{x_{*}}$ centered at $x_{*}$ with $l(I_{x_{*}})=l(I_{k^{\prime}}),r(I_{x_{*}})=r(I_{k^{\prime\prime}})$ for some $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$.
@@
Assume $\{\rho_{n}\}$ contains a subsequence converging to 00. Then $Mf(x)=\frac{f(x^{-})+f(x^{+})}{2}$. This implies $f(x^{-})=f(x^{+})=Mf(x)$ by $Mf(x)\geq Mf(x^{\pm})\geq f(x^{\pm})$, where the first inequality holds because $x$ is a local maximum. Note that $x\leq A$ cannot happen because then $Mf(x)=f(x^{-})=a$ so that $Mf(y)\geq Mf(x)$ for all $y\leq x$. Similarly, $x\geq B$ is impossible. Thus, $x\in I_{k}$ and $Mf(x)=\alpha_{k}$ for some $k$. We can replace $x$ by $c(I_{k})$, since $Mf(y)\geq Mf(x)$ for all $y\in I_{k}$ so that $Mf$ is constant on $I_{k}$ because $x$ is a local maximum. Taking $x_{*}=c(I_{k})$ and $I_{x_{*}}=I_{k}$, we obtain the claimed property.
@@
If $\{\rho_{n}\}$ is separated from 00, then it contains a subsequence converging to a positive number. Consequently, $Mf(x)=\frac{1}{|I_{x}|}\int_{I_{x}}f$ for some $I_{x}$ centered at $x$. Note that $f\big{(}l(I_{x})^{+}\big{)}\geq Mf(x)$, since otherwise we could show $Mf(y)>Mf(x)$ for some $\delta>0$ and all $y\in(x,x+\delta)$ by taking $I_{y}$ with $r(I_{y})=r(I_{x})$. Moreover, $l(I_{x})\geq A$ because otherwise we would have $Mf(y)\geq Mf(x)$ for all $y\leq x$. Thus, $l(I_{x})\in\big{[}l(I_{k^{\prime}}),r(I_{k^{\prime}})\big{)}$ for some $k^{\prime}$. Similarly, $f\big{(}r(I_{x})^{-}\big{)}\geq Mf(x)$ and $r(I_{x})\in\big{(}l(I_{k^{\prime\prime}}),r(I_{k^{\prime\prime}})\big{]}$ for some $k^{\prime\prime}\geq k^{\prime}$. Let $x_{*}=\frac{l(I_{k^{\prime}})+r(I_{k^{\prime\prime}})}{2}$. If $x=x_{*}$, then $Mf(x)$ is attained for $\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$, as desired. If not, then we can show $Mf(y)\geq Mf(x)$ for all $y$ in the closed interval connecting $x$ and $x_{*}$ by taking all $I$ with $l(I_{k^{\prime}})\leq l(I)\leq l(I_{x})$ and $r(I_{x})\leq r(I)\leq r(I_{k^{\prime\prime}})$. Since $x$ is a local maximum, $Mf$ is constant on this closed interval and $Mf(x_{*})$ is attained for $\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$, as desired. ∎
@@
We conclude with noting that each substantial local maximum $x$ of $Mf$, if exists, is of the form $x=\frac{l(I_{k^{\prime}})+r(I_{k^{\prime\prime}})}{2}$ for some $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$, and $Mf(x)=\frac{1}{|I_{x}|}\int_{I_{x}}f$ for $I_{x}=\big{(}l(I_{k^{\prime}},r(I_{k^{\prime\prime}})\big{)}$. Moreover, if $Mf$ has local maxima and there are $N\geq 1$ substantial among them, say $x_{1}<x_{2}<\dots<x_{N}$, then for $N\geq 2$ and every $2\leq n\leq N$ we choose a local minimum $y_{n}\in(x_{n-1},x_{n})$. Then
${\rm Var}(Mf)={\rm Var}_{(-\infty,x_{1}]}(Mf)+\Big{(}\sum_{n=2}^{N}Mf(x_{n-1})-2Mf(y_{n})+Mf(x_{n})\Big{)}+{\rm Var}_{[x_{N},\infty)}(Mf),$
with the middle term disappearing when $N=1$.
@@
Let $T_{I}(x)\coloneqq\frac{|I|}{|J_{I}(x)|}$ for all $x\in\mathbb{R}$, where $J_{I}(x)$ is the smallest interval $J\supseteq I$ with $c(J)=x$. The next result says that the variation of $Mf$ is controlled locally by the variation of $T_{I_{k}}$, $1\leq k\leq K$.
@@
Suppose $Mf$ has $N\geq 2$ substantial local maxima $x_{1}<\dots<x_{N}$. Then

### v3

# Variation of the one-dimensional centered maximal operator on simple functions with gaps between pieces

@@ 1. Introduction
It was an important observation of Aldaz and Pérez Lázaro [1] that maximal operators can actually improve the regularity of involved functions rather than simply preserve it. They also pointed out the role to be played by the variation when measuring the regularity. In particular, it was proved in [1] that if $f\in{\rm BV}(\mathbb{R})$, then for the uncentered Hardy–Littlewood maximal operator $\widetilde{M}$ the function $\widetilde{M}f$ is absolutely continuous on $\mathbb{R}$ and ${\rm Var}(\widetilde{M}f)\leq{\rm Var}(f)$. For the centered operator $M$, on the other hand, Kurka [12] proved
@@
with some, quite large, $C>0$. It was conjectured that also in this case $C=1$ is enough, see e.g. [12, Section 1]. In the discrete case, i.e., for $f\in{\rm BV}(\mathbb{Z})$ and the discrete counterpart of $M$, the same was conjectured, see [3, Question B]. It is worth recalling the common truth that the uncentered maximal operator has better regularity behavior than its centered counterpart. Furthermore, the proofs of analogous results are usually much subtler in the centered setting.
It seems that both conjectures, although reasonable and expected to be true, do not take the full advantage of how the limits $f(\pm\infty)$ or $F(\pm\infty)$ for $f\in{\rm BV}(\mathbb{R})$ or $F\in{\rm BV}(\mathbb{Z})$, respectively, determine the limits $Mf(\pm\infty)$ or $MF(\pm\infty)$; see Section 3. We thus propose the following strengthening, which would complement analogous inequalities for other maximal operators; see Remark 4.1.
@@
Our first main result, Theorem 1.1, supports this conjecture. We mention that there is no need to formulate separately the discrete counterpart of Conjecture $\Delta$, since by our second main result, Theorem 2.2, inequality (1.1) and its discrete analogue are equivalent. Clearly, verification of either version of the conjecture, continuous or discrete, reduces to the consideration of nonnegative functions. We shall prove that a certain further reduction to nonnegative simple functions is available; see Proposition 2.3. Finally, we show that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal; see Section 4.
Let $\mathbb{N}\coloneqq\{1,2,\dots\}$. Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{Z}$. For $g\colon\mathbb{K}\to\mathbb{C}$ and any nonempty $E\subseteq\mathbb{K}$ define
@@
for any locally integrable $g\colon\mathbb{R}\to\mathbb{C}$ and $x\in\mathbb{R}$, and for any $G\colon\mathbb{Z}\to\mathbb{C}$ and $n\in\mathbb{Z}$. Of course, if $g\in{\rm BV}(\mathbb{R})$, then $g$ is bounded, and hence so is $Mg$. The same is true for $MG$ when $G\in{\rm BV}(\mathbb{Z})$.
@@
Let $-\infty<A<B<+\infty$, $a,b\in\mathbb{C}$, $K\in\mathbb{N}$, and $\{\alpha_{k}\}_{1}^{K}\subset\mathbb{C}$ such that $\alpha_{k}\alpha_{k+1}=0$, $11\leq k\leq K-$ 1, be given. For a system $-\infty<A=t_{0}<t_{1}<\dots<t_{K}=B<+\infty$ define
$f(x)\coloneqq a{\mathds{1}_{{(-\infty,A)}}}(x)+\sum_{k=1}^{K}\alpha_{k}{\mathds{1}_{{[t_{k-1},t_{k})}}}(x)+b{\mathds{1}_{{[B,\infty)}}}(x)$
for all $x\in\mathbb{R}$. Then
@@
Theorem 1.1 implies (1.2) or (1.3) for all indicator functions of bounded variation, $f\colon\mathbb{R}\to\{0,1\}$ or $F\colon\mathbb{Z}\to\{0,1\}$, respectively, thus strengthening [4, Theorems 1.1 and 1.3]. On the other hand, Theorem 1.1 does not recover [4, Theorem 1.2], which asserts that ${\rm Var}(Mf)\leq{\rm Var}(f)$ holds for every $f\colon\mathbb{R}\to[0,\infty)$ satisfying $f(x)=0$ or $f(x)=Mf(x)$ for almost every $x\in\mathbb{R}$.
Notice the separation condition $\alpha_{k}\alpha_{k+1}=0$ encoded in the systems $\{t_{k}\}_{0}^{K}$ and $\{n_{k}\}_{0}^{K}$. In fact, our techniques allow us to skip this condition at the expense of using $|\alpha_{k}|+|\alpha_{k+1}|$, not $|\alpha_{k+1}-\alpha_{k}|$, on the right-hand side of (1.2) or (1.3). For example, if $\mathbb{K}=\mathbb{Z}$ and $a=b=0$, then (1.3) would take the form ${\rm Var}(MF)\leq 2(|\alpha_{1}|+\dots+|\alpha_{K}|)$. Building on this, one can recover [14, Theorem 1.1], which asserts that ${\rm Var}(MF)\leq 2\|F\|_{\ell^{1}(\mathbb{Z})}$ holds for all $F\in\ell^{1}(\mathbb{Z})$.
Moreover, by combining Theorem 1.1 with [16, Subsection 5.3], we obtain a variant of (1.2) with $M$ replaced by $M^{\theta}$ for any $\theta>0$, where $M^{\theta}$ is the nontangential maximal operator studied in [16].
Lastly, our methods differ from [4] and align more closely with [14]. In particular, the approach in Lemma 3.2 resembles [14, Section 2], though it gains additional efficiency through Lemma 3.1.
Throughout the paper, $I$ with possible subscripts and/or superscripts affixed will denote an open bounded subinterval of $\mathbb{R}$. Then $c(I),l(I),r(I)$ will be the center and the left/right endpoint of $I$, respectively. As already mentioned, ${\rm Var}$ with no subscripts, $\mathbb{R}$ or $\mathbb{Z}$, will denote the total variation; it will be clear from the context to which situation this symbol refers to. Appearance of a subscript will indicate that the variation is related to a subset represented by this subscript.
@@ 2. Transference
$M_{<\delta}\tilde{g}(y)=\max\big{\{}g(x^{+}),(2\delta)^{-1}\big{(}(\delta-y+x)g(x^{-})+(\delta+y-x)g(x^{+})\big{)}\big{\}},$
@@
$|M_{\geq\delta}\tilde{g}(z)-M_{\geq\delta}\tilde{g}(y)|\leq(2\delta)^{-1}(z-y){\rm Var}(g),$
@@
Thus, $M\tilde{g}(x^{+})$ exists and so $\limsup_{t\to x^{+}}Mg(t)-\liminf_{t\to x^{+}}Mg(t)\leq 2\varepsilon$. Letting $\varepsilon\to 0$, we obtain that $Mg(x^{+})$ exists as well. Analogously we verify that $Mg(x^{-})$ exists.
Regarding (2.1), we have $Mg(x)\leq\min\{Mg(x^{-}),Mg(x^{+})\}$ by lower semicontinuity of $Mg$. To prove the opposite inequality, it suffices to consider the case $\min\{Mg(x^{-}),Mg(x^{+})\}>\frac{g(x^{-})+g(x^{+})}{2}$ because $Mg(x)\geq\frac{g(x^{-})+g(x^{+})}{2}.$ Assume $g(x^{-})\leq g(x^{+})$ (the case $g(x^{-})\geq g(x^{+})$ is symmetric). Now $Mg(x^{-})>\max\{g(x^{-}),\frac{g(x^{-})+g(x^{+})}{2}\}$ and so $Mg(y)=M_{\geq\delta}g(y)$ for some $\delta>0$ and all $y\in(x-\delta,x)$. Since, as before, $|M_{\geq\delta}g(x)-M_{\geq\delta}g(y)|\leq(2\delta)^{-1}(x-y){\rm Var}(g)$ for all such $y$, the proof is complete. ∎
@@
Then ${\rm Var}(g)={\rm Var}(G)$ and so $g\in{\rm BV}_{a}^{b}(\mathbb{R})$. Note that $Mg(n)=MG(n)$, $n\in\mathbb{Z}$, since, for $m\in\mathbb{N}$, one of the averages of $|g|$ over the centered intervals $I$ of lengths $|I|=2m\pm 1$ dominates those with $|I|\in[2m-1,2m+1]$. Thus, ${\rm Var}(MG)={\rm Var}_{\mathbb{Z}}(Mg)\leq{\rm Var}(Mg)$, and the claimed inequalities hold.
In the opposite direction, we claim that $C(\mathbb{Z})\geq C(\mathbb{R})$ and $c_{a}^{b}(\mathbb{Z})\leq c_{a}^{b}(\mathbb{R})$, and use a convenient variant of a sampling procedure. Obviously, we can assume $C(\mathbb{Z})<\infty$. Given $g\in{\rm BV}_{a}^{b}(\mathbb{R})$ and $\varepsilon>0$, in view of (2.1), there is $N_{*}=N_{*}(\varepsilon)\in\mathbb{N}$ such that for $E_{N_{*}}\coloneqq 2^{-N_{*}}\mathbb{Z}\cap[-2^{N_{*}},2^{N_{*}}]$ we have
@@
so that $\tilde{G}_{N}\in{\rm BV}_{|a|}^{|b|}(\mathbb{Z})$ and $G_{N}\in{\rm BV}_{a}^{b}(\mathbb{Z})$. Also, $MG_{N}(n)\leq M\tilde{G}_{N}(n)=M_{N}g(n/2^{N})$ for $n\in\mathbb{N}$.
Now, we show that $MG_{N}(n)\geq M\tilde{G}_{N}(n)-\varepsilon/2^{2N_{*}+4}$ for all $n\in E_{N_{*}}^{N}\coloneqq\{2^{N}Q:Q\in E_{N_{*}}\}\subset\mathbb{N}$, provided that $N\geq\max\{N_{*},N_{0}\}$ is large enough. Given $Q\in E_{N_{*}}$, we have
$|I_{Q}|^{-1}\int_{I_{Q}}|g|\geq M\tilde{G}_{N}(2^{N}Q)-\varepsilon/2^{2N_{*}+5}$
for some $I_{Q}$ centered at $Q$. For each $N$ large enough, let $I_{Q}^{N}$ be the longest interval centered at $Q$, contained in $I_{Q}$, with $|I_{Q}^{N}|\in(2\mathbb{N}-1)/2^{N}$. Denoting $g_{N}\coloneqq\sum_{n\in\mathbb{Z}}G_{N}(n){\mathds{1}_{{[(n-1/2)/2^{N},(n+1/2)/2^{N})}}}$, we see that $\lim_{N\to\infty}|I_{Q}^{N}|^{-1}\int_{I_{Q}^{N}}|g_{N}|=|I_{Q}|^{-1}\int_{I_{Q}}|g|$ by the dominated convergence theorem on $I_{Q}$, because $g$ is continuous almost everywhere and bounded. Since $MG_{N}(2^{N}Q)\geq|I_{Q}^{N}|^{-1}\int_{I_{Q}^{N}}|g_{N}|$ and $E_{N_{*}}$ is finite, $MG_{N}(n)\geq M\tilde{G}_{N}(n)-\varepsilon/2^{2N_{*}+4}$ for some $N\geq\max\{N_{*},N_{0}\}$, as desired. Thus,
@@
which contradicts our assumption $C(\mathbb{Z})<\infty$. Consequently, ${\rm Var}(Mg)<\infty$ (this includes the case $C(\mathbb{R})<\infty$). Applying successively the first part of (2.2), (2.4), the definitions of $C(\mathbb{Z})$ and $c_{a}^{b}(\mathbb{Z})$ in the context of $G_{N}\in{\rm BV}_{a}^{b}(\mathbb{Z})$, and (2.3) for $G_{N}$, and then letting $\varepsilon\to 0$, we verify the claim. ∎
@@
(3.1) $\max\{Mg(y^{\prime}),Mg(y^{\prime\prime})\}<\sup\{Mg(y):y\in\mathbb{K}\cap[y^{\prime},y^{\prime\prime}]\}=Mg(x);$
@@
Until the end of this section we fix $f$ as in Theorem 1.1 and denote $I_{k}\coloneqq(t_{k-1},t_{k})$, $k=1,\ldots,K$, the interiors of the corresponding intervals. Without any loss of generality we may consider only nonnegative $f$. Hence, from now on, we assume $a,b\geq 0$ and $\alpha_{k}\geq 0$, $k=1,\ldots,K$. Clearly, $Mf$ may not have local maxima at all. However, if $Mf$ has a local maximum at $x$, then the following two cases may occur:
@@
    $Mf$ is constant on some interval $J\ni x$, i.e., $Mf$ has a local maximum at each point of $J$;
@@
In the first case, considering $J$ as the longest interval with the declared property, we shall choose exactly one element $x_{*}\in J$ as a representative. In addition every isolated point $x$ from the second case, if any, will also be called a representative. It follows from Lemma 3.1 that the number of representatives is finite and does not exceed $K^{2}$.
@@
Let $Mf$ have a local maximum at $x$. Then there exists $x_{*}$ (either equal to $x$ or such that $Mf$ is constant on the closed interval connecting $x$ and $x_{*}$) with the property that $Mf(x_{*})$ is attained for some interval $I_{x_{*}}$ which is centered at $x_{*}$ and such that $l(I_{x_{*}})=l(I_{k^{\prime}})$ and $r(I_{x_{*}})=r(I_{k^{\prime\prime}})$ hold for some $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$ satisfying $Mf(x_{*})\leq\min\{\alpha_{k^{\prime}},\alpha_{k^{\prime\prime}}\}$.
@@
Assume $\{\rho_{n}\}$ has a subsequence converging to 00. Then $Mf(x)=\frac{f(x^{-})+f(x^{+})}{2}$ and so $f(x^{-})=f(x^{+})=Mf(x)$ by $Mf(x)\geq Mf(x^{\pm})\geq f(x^{\pm})$, where the first inequality holds because $Mf$ has a local maximum at $x$. If $x\leq A$, then $Mf(x)=f(x^{-})=a=\frac{f(y^{-})+f(y^{+})}{2}\leq Mf(y)$ for all $y<x$, contradicting (3.1). Similarly, $x\geq B$ is impossible. Thus, $x\in I_{k}$ and $Mf(x)=\alpha_{k}$ for some $k$. We can replace $x$ by $c(I_{k})$, since $Mf(y)\geq Mf(x)$ for all $y\in I_{k}$ so that $Mf$ is constant on $I_{k}$ because $Mf$ has a local maximum at $x$. Taking $x_{*}=c(I_{k})$ and $I_{x_{*}}=I_{k}$, we obtain the claimed property.
@@
If $\{\rho_{n}\}$ is separated from 00, then it contains a subsequence converging to a positive number. Consequently, $Mf(x)=\frac{1}{|I_{x}|}\int_{I_{x}}f$ for some $I_{x}$ centered at $x$. Note that $f\big{(}l(I_{x})^{+}\big{)}\geq Mf(x)$, since otherwise we could show $Mf(y)>Mf(x)$ for some $\delta>0$ and all $y\in(x,x+\delta)$ by taking $I_{y}$ with $r(I_{y})=r(I_{x})$. In particular, $l(I_{x})\geq A$ because otherwise we would have $Mf(y)\geq Mf(x)$ for all $y\leq x$, again by taking $I_{y}$ with $r(I_{y})=r(I_{x})$. Thus, $l(I_{x})\in\big{[}l(I_{k^{\prime}}),r(I_{k^{\prime}})\big{)}$ for some $k^{\prime}$ such that $\alpha_{k^{\prime}}\geq Mf(x)$. Similarly, $f\big{(}r(I_{x})^{-}\big{)}\geq Mf(x)$ and $r(I_{x})\in\big{(}l(I_{k^{\prime\prime}}),r(I_{k^{\prime\prime}})\big{]}$ for some $k^{\prime\prime}\geq k^{\prime}$ such that $\alpha_{k^{\prime\prime}}\geq Mf(x)$. Let $x_{*}=\frac{l(I_{k^{\prime}})+r(I_{k^{\prime\prime}})}{2}$. If $x=x_{*}$, then $Mf(x)$ is attained for $\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$. If not, then we can show $Mf(y)\geq Mf(x)$ for all $y$ in the closed interval connecting $x$ and $x_{*}$ by taking $x_{\theta}=(1-\theta)x+\theta x_{*}$ and $I_{x}\subseteq I_{x_{\theta}}\subseteq\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$ with $l(I_{x_{\theta}})=(1-\theta)l(I_{x})+\theta l(I_{k^{\prime}})$ and $r(I_{x_{\theta}})=(1-\theta)r(I_{x})+\theta r(I_{k^{\prime\prime}})$ for all $\theta\in[0,1]$. Since $Mf$ has a local maximum at $x$, it is constant on this closed interval and so $Mf(x_{*})$ is attained for $\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$, as desired.
From the reasoning above, it follows that $Mf(x_{*})\leq\min\{\alpha_{k^{\prime}},\alpha_{k^{\prime\prime}}\}$ in each case. ∎
@@
From now on, we take each representative $x$ to be of the form $\frac{l(I_{k^{\prime}})+r(I_{k^{\prime\prime}})}{2}$ for $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$ such that $Mf(x)=\frac{1}{|I_{x}|}\int_{I_{x}}f$, where $I_{x}=\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$. Furthermore, if $Mf$ has local maxima and there are $N\geq 1$ representatives, say $x_{1}<x_{2}<\dots<x_{N}$, then for $N\geq 2$ and each $2\leq n\leq N$, we choose $y_{n}\in(x_{n-1},x_{n})$, where $Mf$ has a local minimum. Then
(3.2) ${\rm Var}(Mf)={\rm Var}_{(-\infty,x_{1}]}(Mf)+\Big{(}\sum_{n=2}^{N}Mf(x_{n-1})-2Mf(y_{n})+Mf(x_{n})\Big{)}+{\rm Var}_{[x_{N},\infty)}(Mf).$
@@
Let $T_{I}(x)\coloneqq\frac{|I|}{|J_{I}(x)|}$ for all $x\in\mathbb{R}$, where $J_{I}(x)$ is the smallest interval $J\supseteq I$ with $c(J)=x$. The next result says that the variation of $Mf$ is controlled locally by the variations of $T_{I_{k}}$, $1\leq k\leq K$.
@@
Suppose $Mf$ has $N\geq 1$ representatives $x_{1}<\dots<x_{N}$. Then